# Strong Lensing Effect and Quasinormal Modes of Oscillations of Black Holes in $\boldsymbol{f(R,T)}$ Gravity Theory

### v1

@@ V Evolution of a scalar perturbation around the black holes
In this section, we study the evolution of a scalar perturbation around our considered BHs’ spacetime for three different cases: $\omega=0$, $\omega=1/3$ and $\omega=-2/3$. We use the time domain integration method to study this evolution. For this, we follow the method described in Refs. [133, 134]. Thus, considering $\psi(x,t)=\psi(i{\Delta}x,j{\Delta}t)=\psi_{i,j}$ and $V(r(x))=V(x,t)=V_{i,j}$, the equation (43) can be expressed in the following form:
@@
Figure 12: Time domain profile for the scalar perturbation in the spacetime of the minimally coupled $f(R,T)$ BHs. The left plot is for $\omega=0$ and $\beta=-8.8$, the right plot is for $\omega=1/3$ and $\beta=-6$, and the bottom plot is for $\omega=-2/3$ and $\beta=-2$.
@@
Figure 13: Fitting of the time domain profile to estimate the QNMs as an example for $l=1$. The left plot is for $\omega=0$, $\beta=-8.8$, the right plot is for $\omega=1/3$, $\beta=-6$ and the bottom plot is for $\omega=-2/3$, $\beta=-2$.
Fig. 13 shows the fitting of the time domain profiles using the Levenberg Marquardt algorithm [135, 136, 137]. From the fitting, we calculate the QNMs of the BHs. Further, we calculate the difference in the magnitudes of the QNMs obtained from the 3rd order WKB approximation method and those obtained from the time domain analysis method as follows:
@@
Table 7 shows the calculated QNMs of the BHs for $\omega=0$ with $\beta=-8.8$, $\omega=1/3$ with $\beta=-6$, and $\omega=-2/3$ with $\beta=-2$ respectively.
Table 7: QNMs of the minimally coupled $f(R,T)$ gravity BHs computed from the fitting of the time domain profiles using the Levenberg Marquardt algorithm [135, 136, 137] in comparison with that obtained from 3rd order WKB approximation method for different values of $\omega$ and $\beta$.
$\omega$ $\beta$ Multipole $l$ 3rd order WKB QMNs Time domain QNMs $R^{2}$ $\Delta_{Q\!N\!M}$
$\omega=0$ $\beta=-8.8$ $1$ $0.371088-0.116790i$ $0.371218-0.118280i$ 0.995280 $2.818530\times 10^{-4}$
$2$ $0.606594-0.098307i$ $0.602417-0.097901i$ 0.985772 $2.093900\times 10^{-3}$
$3$ $0.846850-0.092873i$ $0.842375-0.091867i$ 0.988405 $2.278800\times 10^{-3}$
$\omega=1/3$ $\beta=-6$ $1$ $0.333450-0.087057i$ $0.324891-0.149260i$ 0.981999 $6.455010\times 10^{-3}$
$2$ $0.554941-0.086645i$ $0.522741-0.095422i$ 0.987519 $1.514250\times 10^{-2}$
$3$ $0.775963-0.086548i$ $0.723183-0.091149i$ 0.989405 $2.593500\times 10^{-2}$
$\omega=-2/3$ $\beta=-2$ $1$ $1.341550-0.494087i$ $1.343261-0.472752i$ 0.971959 $2.808810\times 10^{-3}$
$2$ $1.972000-0.537327i$ $1.971946-0.516896i$ 0.989519 $2.664080\times 10^{-3}$
$3$ $2.635360-0.554069i$ $2.635490-0.554137i$ 0.989134 $7.061500\times 10^{-5}$
@@ VI Conclusion
Gravitational lensing is a versatile technique for exploring the properties of spacetime and for testing the predictions of different theories of gravity. In the first part of this study, we have analyzed the strong gravitational lensing properties of BHs described by the $f(R,T)$ gravity theory, focusing on a minimally coupled model of the theory. The analysis has relied on a strong field limit universal method developed by V. Bozza, applicable to all spacetime in any gravity theory where the photons follow the standard geodesic equations [43]. We have conducted the analysis considering three types of surrounding fields around the BHs, namely the dust field, the radiation field and the quintessence field. Here, we have concentrated on the investigation of the characteristics of the region just outside the event horizon of the minimally coupled $f(R,T)$ gravity BHs for these fields separately through the analysis of photon deflection extremely close to the photon sphere of the BHs. For this, the influence of the model parameter $\beta$ on different photon trajectories is examined. Our analysis reveals that the lensing coefficients $\bar{a}$ and $\bar{b}$ exhibit distinct and contrasting behaviors with $\beta$ depending on the values of $\omega$. These variations directly affect the deflection angle $\hat{\alpha}$ which decreases with increasing impact parameter $\zeta$ and becomes negative for larger values of $\zeta$, highlighting repulsive interactions by the gravitational field of the BHs. Furthermore, by considering the supermassive BH SgrA* as a minimally coupled $f(R,T)$ gravity BH, we have computed the numerical values of the lensing observables. The computed lensing observables: the angular position of relativistic images $\vartheta_{\infty}$, angular separation $s$, and relative magnification $r_{\text{mag}}$ demonstrate a consistent decline with increasing $\beta$ in scenarios with $\omega=0$ and $\omega=1/3$. Remarkably, when comparing the observables of the first two classes of BHs, the respective observables are found smaller in a dust dominated environment than in the presence of the radiation field, except for $\beta\leq-12.23$, in which a sudden growth of dust field observables is noticed. This sudden increase of observables could be due to the dominance of the physical parameters linked to the dust field over those of the radiation field beyond this value. In addition, the analysis for $\omega=-2/3$ reveals a distinct deviation in the behavior of the observables beyond $\beta\approx-9.8$ in contrast to the cases of $\omega=0$ and $\omega=1/3$. The sudden decline of $\vartheta_{\infty}$ and the steep growth of $s$ and $r_{mag}$ up to $\beta=-9.4381$ indicate a significant dependence of the observables on the parameter $\beta$. These results highlight that reliable predictions for the observables are confined within this specific range of $\beta$, emphasizing the intricate relationship between $\beta$ and the underlying physical properties of the system. However, our findings underscore the significant role of $\beta$ and the type of surrounding field in shaping the strong lensing behavior of the considered BHs spacetime with deviation limits $\beta=-16.7551$ in $\omega=0$ case, $\beta=-18.8495$ in $\omega=1/3$ and $\beta=-13.7088$ for $\omega=-2/3$, from GR. Indeed, prominent deviation is observed as $\beta$ increases providing a potential avenue for testing deviations from GR through strong lensing observations.
@@
The QNMs associated with the BHs’ spacetime have also been studied. For three values of equation of state parameter $\omega$, we calculated the QNMs of BHs using the 3rd order WKB approximation method as the higher order ($>$ 5th order) values of QNMs of the BHs obtained by this method are not converging. We plot the real part of calculated QNMs denoting the amplitude of frequency of the mode against the model parameter $\beta$. It is seen that the amplitude increases with increasing $\beta$ for all three cases. For the imaginary part of the QNMs denoting damping rate, we found that for $\omega=-2/3$, the highest damping has been observed compared to the other two cases. For $\omega=1/3$, the plots converse near $\beta=0$ and then show a steady increase in damping as $\beta$ increases. Time evolution of the QNMs of oscillations has also been studied and the decaying modes are clearly visible from their profiles. For the sake of accuracy, we have compared the results obtained by the WKB method with that obtained from the time profile fitting method based on the Levenberg Marquardt algorithm and found a good agreement in results as shown in Table 7. The deviation between the results of these two methods has been represented via a quantity parameter $\Delta_{Q\!N\!M}$ defined in equation (54), which comes out to be of the order of about $10^{-3}$, thus signifying a good match between the two methods.

### v2

@@ V Evolution of a scalar perturbation around the black holes
In this section, we study the evolution of a scalar perturbation around our considered BHs’ spacetime for three different cases: $\omega=0$, $\omega=1/3$ and $\omega=-2/3$ for different values of $\beta$ and $c_{2}$. We use the time domain integration method to study this evolution. For this, we follow the method described in Refs. [153, 154]. Thus, considering $\psi(x,t)=\psi(i{\Delta}x,j{\Delta}t)=\psi_{i,j}$ and $V(r(x))=V(x,t)=V_{i,j}$, the equation (43) can be expressed in the following form:
@@
Refer to caption   Refer to caption    Refer to caption
Figure 18: Time domain profile for the scalar perturbation in the spacetime of the minimally coupled $f(R,T)$ BHs. In the upper row, the left plot is for $\omega=0$, $\beta=-8.8$ and $c_{2}=1$, the middle plot is for $\omega=1/3$, $\beta=-6$ and $c_{2}=1$, and the right plot is for $\omega=-2/3$, $\beta=-2$ and $c_{2}=1$. In the lower row, the left plot is for $\omega=0$, $\beta=-8.8$ and $c_{2}=0.1$, the middle plot is for $\omega=1/3$, $\beta=-5.5$ and $c_{2}=0.5$, and the right plot is for $\omega=-2/3$, $\beta=-2$ and $c_{2}=0.8$.
@@
Refer to caption    Refer to caption   Refer to caption
Figure 19: Fitting of the time domain profile to estimate the QNMs as an example for $l=1$. In the upper row, the left plot is for $\omega=0$, $\beta=-8.8$ and $c_{2}=1$, the middle plot is for $\omega=1/3$, $\beta=-6$ and $c_{2}=1$, and the right plot is for $\omega=-2/3$, $\beta=-2$ and $c_{2}=1$. In the lower row, the left plot is for $\omega=0$, $\beta=-8.8$ and $c_{2}=1.1$, the middle plot is for $\omega=1/3$, $\beta=-5.5$ and $c_{2}=0.5$, and the right plot is for $\omega=-2/3$, $\beta=-2$ and $c_{2}=0.8$.
Fig. 19 shows the fitting of the time domain profiles using the Levenberg Marquardt algorithm [155, 156, 157]. From the fitting, we calculate the QNMs of the BHs. Further, we calculated the difference in the magnitudes of the QNMs obtained from the 3rd order WKB approximation method and those obtained from the time domain analysis method as follows:
@@
Table 6 shows the calculated QNMs of the BHs for $\omega=0$, $\omega=1/3$ and $\omega=-2/3$ with different values of $\beta$ and $c_{2}$.
Table 6: QNMs of the minimally coupled $f(R,T)$ gravity BHs computed from the fitting of the time domain profiles using the Levenberg Marquardt algorithm [155, 156, 157] in comparison with that obtained from 3rd order WKB approximation method for different values of $\omega$, $\beta$, and $c_{2}$.
$\omega$ $\beta$ $c_{2}$ Multipole $l$ 3rd order WKB QMNs Time domain QNMs $R^{2}$ $\Delta_{Q\!N\!M}$
$\omega=0$ $\beta=-8.8$ $c_{2}=1$ 11 0.3710880.116790i0.371088-0.116790i 0.3712180.118280i0.371218-0.118280i 0.995280 $2.818530\times 10^{-4}$
22 0.6065940.098307i0.606594-0.098307i 0.6024170.097901i0.602417-0.097901i 0.985772 $2.093900\times 10^{-3}$
33 0.8468500.092873i0.846850-0.092873i 0.8423750.091867i0.842375-0.091867i 0.988405 $2.278800\times 10^{-3}$
$\omega=1/3$ $\beta=-6$ $c_{2}=1$ 11 0.3334500.087057i0.333450-0.087057i 0.3248910.149260i0.324891-0.149260i 0.981999 $6.455010\times 10^{-3}$
22 0.5549410.086645i0.554941-0.086645i 0.5227410.095422i0.522741-0.095422i 0.987519 $1.514250\times 10^{-2}$
33 0.7759630.086548i0.775963-0.086548i 0.7231830.091149i0.723183-0.091149i 0.989405 $2.593500\times 10^{-2}$
$\omega=-2/3$ $\beta=-2$ $c_{2}=1$ 11 1.3415500.494087i1.341550-0.494087i 1.3432610.472752i1.343261-0.472752i 0.971959 $2.808810\times 10^{-3}$
22 1.9720000.537327i1.972000-0.537327i 1.9719460.516896i1.971946-0.516896i 0.989519 $2.664080\times 10^{-3}$
33 2.6353600.554069i2.635360-0.554069i 2.6354900.554137i2.635490-0.554137i 0.989134 $7.061500\times 10^{-5}$
$\omega=0$ $\beta=-8.8$ $c_{2}=0.1$ 11 0.3097610.123334i0.309761-0.123334i 0.3095430.122674i0.309543-0.122674i 0.991752 $3.818000\times 10^{-4}$
22 0.4924410.105570i0.492441-0.105570i 0.4936730.106510i0.493673-0.106510i 0.997334 $0.775000\times 10^{-4}$
33 0.6862220.101014i0.686222-0.101014i 0.6851220.102673i0.685122-0.102673i 0.985176 $0.997300\times 10^{-3}$
$\omega=1/3$ $\beta=-5.5$ $c_{2}=0.5$ 11 0.3191040.122256i0.319104-0.122256i 0.3111720.122400i0.311172-0.122400i 0.993911 $3.967500\times 10^{-3}$
22 0.5156560.104737i0.515656-0.104737i 0.5157230.103974i0.515723-0.103974i 0.99632 $4.283255\times 10^{-5}$
33 0.7184150.0998847i0.718415-0.0998847i 0.7182430.093451i0.718243-0.093451i 0.985176 $3.217500\times 10^{-3}$
$\omega=-2/3$ $\beta=-3$ $c_{2}=0.8$ 11 1.1494900.424259i1.149490-0.424259i 1.1276320.423921i1.127632-0.423921i 0.981517 $1.093100\times 10^{-2}$
22 1.7119800.455701i1.711980-0.455701i 1.7098500.454871i1.709850-0.454871i 0.98532 $1.343781\times 10^{-3}$
33 2.3007600.467673i2.300760-0.467673i 2.3546100.461734i2.354610-0.461734i 0.985176 $2.708001\times 10^{-2}$
@@ VI Conclusion
Gravitational lensing is a versatile technique for exploring the properties of spacetime and for testing the predictions of different theories of gravity. In the first part of this study, we have analyzed the strong gravitational lensing properties of BHs described by the $f(R,T)$ gravity theory, focusing on a minimally coupled model of the theory. The analysis has relied on a strong field limit universal method developed by V. Bozza, applicable to all spacetime in any gravity theory where the photons follow the standard geodesic equations [48]. We have conducted the analysis considering three types of surrounding fields around the BHs, namely the dust field, the radiation field and the quintessence field. Here, we have concentrated on the investigation of the characteristics of the region just outside the event horizon of the minimally coupled $f(R,T)$ gravity BHs for these fields separately through the analysis of photon deflection extremely close to the photon sphere of the BHs. For this, the influences of the model parameters $\beta$ and $c_{2}$ on different photon trajectories are examined. Our analysis reveals that the lensing coefficients $\bar{a}$ and $\bar{b}$ exhibit distinct and contrasting behaviors with $\beta$ depending on the values of $\omega$, and result in nearly identical trends for both $\omega=0$ and $1/3$, however, show contrasting behavior for $\omega=-2/3$ with $c_{2}$. These variations directly affect the deflection angle $\hat{\alpha}$ which decreases with increasing impact parameter $\zeta$ and becomes negative for larger values of $\zeta$, highlighting repulsive interactions by the gravitational field of the BHs. Furthermore, by considering the supermassive BH SgrA* as a minimally coupled $f(R,T)$ gravity BH, we have computed the numerical values of the lensing observables. The computed lensing observables: the angular position of relativistic images $\vartheta_{\infty}$, angular separation $s$, and relative magnification $r_{\text{mag}}$ demonstrate a consistent decline with increasing $\beta$ and $c_{2}$ in scenarios with $\omega=0$ and $\omega=1/3$. Remarkably, when comparing the observables of the first two classes of BHs, the respective observables are found smaller in a dust dominated environment than in the presence of the radiation field. Eventually, the variation as a function of $\beta$ shows a sudden growth of dust field observables beyond $\beta\leq-12.23$. This sudden increase of observables could be due to the dominance of the physical parameter linked to the dust field over those of the radiation field beyond this value. In addition, the analysis for $\omega=-2/3$ reveals a distinct deviation in the behavior of the observables beyond $\beta\approx-9.8$ in contrast to the cases of $\omega=0$ and $\omega=1/3$. The sudden decline of $\vartheta_{\infty}$ and the steep growth of $s$ and $r_{mag}$ beyond $\beta=-9.8$ indicate a significant dependence of the observables on the parameter $\beta$. These results highlight that reliable predictions for the observables are confined within this specific range of $\beta$, emphasizing the intricate relationship between $\beta$ and the underlying physical properties of the system. However, our findings underscore the significant role of $\beta$ and the type of surrounding field in shaping the strong lensing behavior of the considered BHs spacetime with deviation from the GR, greater in case of $\omega=0$, and lesser in $\omega=-2/3$ situation respectively for any positive value of $c_{2}$. Additionally, we have constrained parameters $\beta$ and $c_{2}$ from the EHT observational data of SgrA* to test the viability and observational consistency of the model. Indeed, the constraining of the parameters enables us to confirm a limit to $\beta$ and $c_{2}$ in such a way that the upper bound of $\beta$ increases with the decrease of $c_{2}$ providing a potential avenue for examining analyses through strong lensing observations.
@@
The QNMs associated with the BHs’ spacetime have also been studied. For three values of equation of state parameter $\omega$, we calculated the QNMs of BHs using the 3rd order WKB approximation method as the higher order (>> 5th order) values of QNMs of the BHs obtained by this method are not converging. We plot the real part of calculated QNMs denoting the amplitude of frequency of the mode against the model parameter $\beta$. It is seen that the amplitude increases with increasing $\beta$ for all three cases. For the imaginary part of the QNMs denoting damping rate, we found that for $\omega=-2/3$, the highest damping has been observed compared to the other two cases. For $\omega=1/3$, the plots converse near $\beta=0$ and then show a steady increase in damping as $\beta$ increases. The analysis is done again for varying $c_{2}$ parameter. It is seen that the parameter has profound impact on the amplitude as well as damping of QNMs, with its influence increasing for higher multipole $l$. Time evolution of the QNMs of oscillations has also been studied and the decaying modes are clearly visible from their profiles. For the sake of accuracy, we have compared the results obtained by the WKB method with that obtained from the time profile fitting method based on the Levenberg Marquardt algorithm and found a good agreement in results as shown in Table 6. The deviation between the results of these two methods has been represented via a quantity parameter $\Delta_{Q\!N\!M}$ defined in equation (54), which comes out to be of the order of about $10^{-3}$, thus signifying a good match between the two methods.